OpenStudy (anonymous):

what does it mean for a number to be the "solution" to an equation? 1.) give an example of a variable equation and its solution help?

OpenStudy (anonymous):

X+5 = 0 ---> equation solve for x, means to find the values of x for which the equation is true. x=-5, and if you plug it in the equation, (-5)+5 = 0, the equation is true, thus x=-5 is the soln

OpenStudy (anonymous):

how would you explain this in complete sentences how yu know that number is the solution?

OpenStudy (anonymous):

A solution for a variable is basically a value of the variable which results in equal right and left hand-sides of an equation. For example, If I am give the equation:\[\bf y-3=2\]A solution to this problem would be a value of 'y' such that when that value is substituted for 'y', the left and right hand-sides of the equation are the same. Here we can re-arrange the equation like this:\[\bf y=2+3=5\]Hence y = 5 is a solution to the equation because plugging in 5 for 'y' in to the original equation gives us the same left and right hand sides.

OpenStudy (anonymous):

thats how i would answer both these questions?

OpenStudy (anonymous):

cuz it says give example of a variable equation and its solution then it says explain in complete sentences how you know that number is the solution